OpenStudy (anonymous):

What formula should I use to solve this? I don't want the answer just the formula. There are six snakes in a valley. The population of snakes doubles every year. In how many years will there be 96 snakes?

OpenStudy (anonymous):

solve \[6\times 2^n=96\] for n

OpenStudy (anonymous):

divide by 6 get \[2^n=16\] so \[n=4\] because \[2^4=16\]

OpenStudy (anonymous):

to solve that, divide both sides by 6 and then take \[\log_2(n)\] of both sides \[n=\log_2(16)=4\log(2) = 4\]